OpenStudy (anonymous):

What is the area, rounded to the nearest tenth, of a circle with a diameter of 22 m? Choose one answer. a. 1520.5 m^2 b. 1194.2 m^2 c. 69.1 m^2 d. 380.1 m^2

OpenStudy (anonymous):

wow thanks

OpenStudy (anonymous):

The area for a circle can be found like so,\[Area = \pi r^2\] where r = radius and 1/2 diameter = radius. \[Area = π× (\frac{ 22 }{ 2 })^{2} = \pi \times 11^{2}≈380.1\]